OpenStudy (anonymous):

cos ( 2cos^-1 x + sin^-1 x) at x= 1/7 = ?

OpenStudy (anonymous):

cos ( cos^-1 1/7+ cos^-1 1/7 + sin^-1 1/7)

OpenStudy (anonymous):

=cos ( cos^-1 1/7 + pie/2)

OpenStudy (anonymous):

It seems wrong (IMHO of course, may be not...)

OpenStudy (anonymous):

cos (cos^ -1 1/7) cos pie/2 - sin (cos^ -1 1/7) sin pie/2

hartnn (hartnn):

cos pi/2 = 0

OpenStudy (shubhamsrg):

sin x = sqrt(1-cos^2 x) ..maybe this helps..

OpenStudy (anonymous):

Is this correct

OpenStudy (anonymous):

0 - sin ( pie/2 - sin^-1 1/7)

hartnn (hartnn):

use this, u will get it sin^-1 t + cos^-1 t =pi/2

OpenStudy (shubhamsrg):

-sin (cos^ -1 1/7) = - sqrt ( 1- cos(cos-1 (1/7) ^2 ) ) = -sqrt ( 1- 1/49) was that right ?

hartnn (hartnn):

how does sqrt come into picture ??

OpenStudy (anonymous):

i didnt Understand hw u reached here

OpenStudy (shubhamsrg):

just to eliminate the inverse term,i used sinx = sqrt(1- cos^2 x) ..

OpenStudy (anonymous):

The answer is correct

OpenStudy (anonymous):

lol....Yes....)...thxxxx

OpenStudy (anonymous):

-4sqrt3/7

OpenStudy (shubhamsrg):

yep..the very same..